Reports 1-1 of 1 Clear search Modify search
PEM (Center)
takahiro.yamamoto - 2:58 Friday 16 April 2021 (16488) Print this report
template for checking survival situation of PEM sensor
[Yokozawa, Yamamoto]

Yokozawa-san restored PEM sensors in PSL room today.
We are not sure how long it has been stopped (maybe from last planned power outage?).

Since we often lose track of the operational status of PEM sensors, we have prepared a diaggui template to quickly check their survival. (See figure1)
In this template, reference spectra after today's recovery and during O3GK.
Because many FFUs are running, current noise level is louder than one during O3GK.

And also we checked the coherence between PEM sensors and calibrated PMC frequency noise.
As same as the situation during O3GK, microphone has coherence in the range of 10-200Hz. (See figure2).

What was interesting was that even before the PEM sensor was turned back on, some channels showed coherence with only circuit and cable noise being seen. ((See figure3)
From my experience, ADC (and AA) noise without connection to other instruments doesn't have such a large cross talk.
So we had better to check the cabling and/or electrical connection around IOO rack and PSL room.
Images attached to this report
Comments to this report:
takaaki.yokozawa - 7:31 Friday 16 April 2021 (16490) Print this report
We want to discuss detail in some meeting, but I want to leave some comments.

K1:PEM-ACC_PSL_PERI_PSL1_Y_OUT is not connected
(No refcav now, no periscope now, no PEM(Where??))

Coherence between PMC control signal and microphone was detected long time,
we want to check the spectrum of origin of sound inside the PSL room.

As Yamamoto-san reported, no PEMs was worked until yesterday.
(Turned off the drivers)
As ACC_TABLE1 and ACC_TABLE3 has some excess compared from ADC noise
And interesting issue is the behavior between ACC_TABLE2 and ACC_TABLE1,ACC_TABLE3 is slightly different(Noise floor and floor shape)
But we couldn't find the reason...
hirotaka.yuzurihara - 17:55 Tuesday 20 April 2021 (16514) Print this report

[Yokozawa, Washimi, Yuzurihara]

Today we entered the PSL room from 14:00~15:00.

To make the concrete plan to find the origin of the coherence between the microphone installed at PSL and PMC feedback signal, I used the portable microphone and the chromebook and I saw the spectrum at each place in PSL room.

The comparisons of the spectrum will be done later.

Search Help